OpenStudy (anonymous):

How many solutions are there to the following system of equations? 3x - 9y=0 -x + 3=-3 A. 0 B. 2 C. 1 D. infinitely many Will medal/fan

OpenStudy (anonymous):

@confluxepic

OpenStudy (anonymous):

@Loser66

OpenStudy (confluxepic):

@TheSmartOne

TheSmartOne (thesmartone):

Lets solve the 2 equations.

TheSmartOne (thesmartone):

So which method do you like to solve them by? Substituion, elimination, or graphing?

OpenStudy (anonymous):

Substituion.

TheSmartOne (thesmartone):

So lets solve for x in this equation \(\sf -x + 3=-3\)

TheSmartOne (thesmartone):

We can add x to both sides and then add 3 on both sides. And we will have isolated x.

OpenStudy (anonymous):

like this?\[3=-3+x?\]

TheSmartOne (thesmartone):

yup and then just add 3 on both sides again :)

OpenStudy (anonymous):

\[3+3=-3+x+3?\]

TheSmartOne (thesmartone):

and \(\sf -3+3 =0\) And what is \(\sf 3+3=?\)

OpenStudy (anonymous):

6

TheSmartOne (thesmartone):

So now we have these 2 equations \(\sf 3x - 9y=0 \\x=6\)

TheSmartOne (thesmartone):

So lets plug in x=6 into 3x - 9y=0 \(\sf 3x - 9y=0\) \(\sf 3\times (6) - 9y=0\) Solve for y :)

TheSmartOne (thesmartone):

Oh and just a tip, 2 equations can NEVER have 2 solutions. So if you ever see 2 as a answer choice for this type of question, you can eliminate it.

OpenStudy (anonymous):

okay \[3x6=38-9y=29y???\]

TheSmartOne (thesmartone):

ok... what is \(\sf 3 \times 6=?\)

OpenStudy (anonymous):

38

TheSmartOne (thesmartone):

Actually no. \(\ 3 \times 6 = 6+6+6=?\)

OpenStudy (anonymous):

18.... Then I must've failed on my other quizes <.<'

TheSmartOne (thesmartone):

correct. and \(\sf 18-9y=0\) Add 9y to both side and then divide by 9. And you get the vaule of y. But we can stop right here. We have gotten that the answer is (x,y) which is (6,y) (which is what we will get when we solve for y) And (6,2) is only one point. So we only have 1 solution.

OpenStudy (anonymous):

okay.

TheSmartOne (thesmartone):

Well I hope that helps :)